OpenStudy (nuccioreggie):

Part A: 6(2x-11) + 15 = 3x +12 12x-66 + 15 = 3x+12 (distribute 6 times 2x and -11) 12x -51 = 3x+12 (combine -66 and 15) 9x-51 = 12 (subtract 3x from both sides) 9x = 63 (add 51 to both sides) x= 7 (divide both sides by 9) Part B: x=7

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

?????????? What is the question?

OpenStudy (whovianchick):

That's my answer to the last question copied and pasted.

OpenStudy (nuccioreggie):

lol sorry wrong thing thats the answer

OpenStudy (ryuga):

6(2x-11) + 15 = 3x +12

OpenStudy (whovianchick):

@Ryuga there's no need to solve it he said the wrong thing XD

OpenStudy (ryuga):

12x-66 + 15 = 3x+12

OpenStudy (ryuga):

12x -51 = 3x+12

OpenStudy (ryuga):

9x-51 = 12

OpenStudy (ryuga):

9x = 63

OpenStudy (ryuga):

x= 7

OpenStudy (ryuga):

so you are correct

OpenStudy (whovianchick):

That was my answer to his last question XDDD